# ALPINE: Unveiling the Planning Capability of Autoregressive Learning in Language Models

### v2

@@ 6 Path-planning in Blocksworld
To further validate the theoretical results in Section 4 and the practicability of the proposed path-finding task, we consider Blocksworld benchmark [VMSK24]. Blocksworld is a scenario consisting of a set of blocks identified by different colors. The blocks are either placed on table or on top of another block and the task is to plan a block manipulation from the given state to the target state.
@@
We formulate Blocksworld as a path-finding task. Here we construct a graph $G_{BW}$ for the case with $4$ blocks, where each node represents a state of the blocks. For example, node 00 refers to the state that “the red block is on top of the blue block, the blue block is on top of the orange block, the orange block is on top of the yellow block, and the yellow block is on the table”. $G_{BW}$ is a directed graph with $73$ nodes, and the adjacency matrix of $G_{BW}$ is presented in Figure 7(a). Specifically, the states corresponding to the first $24$ nodes are the states where the four blocks are stacked one on top of another in a single stack, and thus each of these nodes only has one out-neighbor corresponding to removing the top block and putting it on the table. The last node refers to the state where all blocks are on the table, so it has 12 out-neighbors, corresponding to the $12$ states where three blocks are on the table and the fourth block is on top of one of the three blocks.
@@
In the original Blocksworld task, the answer is a sequence of actions, which is equivalent to the notion of edges in $G_{BW}$. We reformulated it to let the model output a path from the given state to the target state, only consisting of the nodes. This can be seen as a simplified version and a pure planning task. We randomly select $80\%$ of all node pairs for training and the rest $20\%$ for testing and generate the $50000$ training data in the same format as introduced in Section 5. We mainly use Transformers with $1$ layer and $1$ head for the convenience of visualization.
@@
Figure 8: Reachability for the experiment on Blocksworld benchmark.
6.1 Results
@@
We first present the accuracy during the training when using different embedding sizes $d\in\{30,60,120\}$. As is shown in Figure 7(d), although a smaller embedding size results in a longer time to converge, all models reach an accuracy near $100\%$ at the end of the training.
@@
Then, we use the same method introduced in Section 5 to visualize the attention map and the $\bm{W}^{M^{\prime}}$ matrix for the model with $d=120$ after the entire iterations. In Figure 7(c), we can see that when predicting the tokens on the path, almost all the attention weights are on the second token which represents for the target node, demonstrating the capability of the model to learn a correct attention. For adjacency matrix, we find that the $\bm{W}^{M^{\prime}}$ matrix in Figure 7(b) almost perfectly aligns to the real adjacency matrix of $G_{BW}$. And the weight gap (average edge weight minus average non-edge weight) for all models keeps increasing in the training process until convergence, as is shown in Figure 7(e).
@@
In addition, we present the results related to reachability matrix in Figure 8. Figure 8(a) shows the observed reachability in the training dataset. Although $G_{BW}$ is fully-connected, some reachability are not observed since we request that all training data has no cycle. Specifically, each of the first $24$ nodes is not observed reachable to any nodes other than itself. To validate whether Transformer has captured this information, we construct $\bm{W}^{V^{\prime}}$ matrix through the same method presented in Section 5. As shown in Figure 8(b), the $\bm{W}^{V^{\prime}}$ matrix has represented darkness for the first $24$ columns, which aligns to the real observed reachability matrix. Furthermore, we plot the average gap between the weight of $\bm{W}^{V^{\prime}}$ on observed reachability and those exist but not observed in Figure 8(c). We find that this gap keeps increasing for all models [^4] Since there does not exist any test pairs that are with degree 2 or more (as defined in Section 5), we do not compare the accuracy between different degrees in Blocksworld..
@@
In summary, our experimental results on the Blocksworld benchmark confirms our theoretical analyses (Theorem 4) and empirical results on the synthetic data, and it at least partially explains the planning capability of the Transformer on the Blocksworld scenario.
7 Related Works
7.1 LLMs for Planning
Several recent studies have empirically evaluated the planning abilities of large language models (LLMs). For instance, CogEval has introduced a variety of planning tasks set in mazes and graphs, ultimately finding no evidence to suggest LLMs grasp the intricacies of the maps or the planning tasks themselves [MHVF+23]. Similarly, another study explored the Blocksworld game, a planning challenge where humans typically achieve success rates above $70\%$, in stark contrast to GPT-3’s mere $5\%$ success rate [VMSK24]. Our paper proposes a novel approach by formulating a class of planning problems as path finding on graphs, applying this model to the blocksworld game and uncovering significant insights, as detailed in Section 6.
Despite these seemingly negative evaluations, LLMs have shown remarkable aptitude in executing real-world planning tasks, creating the field of autonomous agents [WMF+24]. Certain applications of autonomous agents feature explicit graphs. In the tool agent HuggingGPT [SST+23], LLMs are deployed to trigger a sequence of external APIs in response to user requests. Here, APIs are conceptualized as graph nodes, with their interrelations represented as edges, and the selection process akin to identifying a path or subgraph that aligns with user demands. This scenario is an extension of the settings discussed in this paper, where the graph is text-attributed and the objective function is evaluated through textual analysis. The application of graph search techniques has been shown to enhance the performance of tool agents significantly [LLG+23, LPY+24]. The math agent AlphaGeometry utilizes LLMs to solve geometry problems [TWL+24]. By treating lemmas as nodes and their interdependencies as edges, the process of finding a proof of a theorem is analogous to finding a path to the theorem node in the above graph formed by possible lemma nodes and their interdependency edges. However, [TWL+24] focuses on using LLM to generate auxiliary constructions, and the reasoning tasks are done by a non-LLM engine. This is very different to our research. There are no explicit graphs in other agents, such as game agents [WXJ+23], embodied agents [HAPM22], and code agents [SCG+24]. The core strategy in these domains is to employ verbal reinforcement learning within LLMs. It is noteworthy that any dynamic programming problem characterized by deterministic state transitions can be reformulated as a shortest path problem on a graph, with states and transitions represented as nodes and edges, respectively. As a result, the area of autonomous agents is closely related to the path-planning task investigated in this paper.
7.2 LLM for Graphs
GPT4Graph [GDL23] and NLGraph [WFH+23] have developed extensive frameworks for assessing LLMs in the context of graph tasks. These frameworks encompass a broad spectrum of challenges, including classic graph problems (e.g., connectivity, cycle detection, and topological sorting), graph neural network (GNN) tasks (e.g., node and graph classification), and semantic graph question answering (e.g., knowledge graph inquiries). They also explore various input formats, such as adjacency lists, edge lists, GML, and GraphML, alongside innovative prompting techniques like few-shot, role prompting, chain-of-thought, and algorithmic prompting (e.g., stating “we are using DFS”). These studies demonstrate that LLMs possess basic graph processing capabilities, and the choice of prompts and formats influence performance significantly. Yet, they also reveal the models’ susceptibility to spurious correlations within graphs. GPT-4, for instance, only achieves around $50\%$ accuracy on shortest path tasks, even when utilizing complex prompts. To our knowledge, this paper presents the first theoretical analysis that identifies and explains the spurious correlations learned by LLMs (Theorem 3), supporting the negative outcomes reported in these studies.
There has also been a surge in efforts aiming at bolstering LLMs’ performance on graph tasks. Innovations like GraphGPT [TYW+23] and GraphLLM [CZW+23], which incorporate an additional GNN encoder, have shown notable improvements across the aforementioned graph tasks. GraphInstruct [LSH+24] seeks to enhance LLMs’ capabilities using pure LLM approaches. This involves meticulously documenting the steps of classical algorithms (e.g., BFS and DFS) and fine-tuning LLMs to learn these graph algorithms. This method of procedural supervision has extended the capacity of LLMs in graph tasks from the complexity class $\text{TC}^{0}$ to P/poly [FZG+23]. However, while this approach has yielded performance improvements in simpler tasks such as topological sorting and connectivity, it has proven less effective for more complex challenges, like finding Hamiltonian Paths.
7.3 Algorithm Simulation with Transformers
Recent theoretical investigations have shed light on the capability of Transformer to simulate algorithms, a topic that has garnered considerable interest. This discussion begins with discrete algorithms. From a circuit complexity standpoint, Transformer models are likened to parallel circuits characterized by polynomial width and constant depth, which places them within the $\text{TC}^{0}$ complexity class. It is also noticed that $\text{TC}^{0}\subseteq\text{NC}^{1}\subseteq P$. On the other hand, despite their impressive expressiveness, Transformer is theoretically incapable of addressing a range of P-complete problems, including the testing of Context-Free Grammar Membership [MS23]. However, the advent of chain-of-thought prompting has enabled Transformer to sequentially simulate algorithms, thereby equipping them to tackle P-complete problems in domains such as arithmetic and decision-making [FZG+23]. The exploration extends to continuous algorithms, where it has been demonstrated that Transformer can approximate functions such as matrix inversion, Stochastic Gradient Descent (SGD), and power iterations [GRS+23]. Our study specifically applies GPT models to simulate path-finding algorithms, presenting evidence that their expressiveness is sufficient for such tasks (Theorem 2). Nevertheless, the usage of auto-regressive loss and gradient descent introduces certain limitations (Theorem 3), which have not been studied in existing works.
7.4 Mechansims of LLMs
LLMs have demonstrated capabilities that exceed the theoretically predicted lower bounds of expressiveness. To demystify this paradox, numerous studies have employed experimental methodologies akin to those used in the physical and biological sciences. Their aim is to decode the mechanisms of LLMs. The foundational strategy is to generate controlled synthetic datasets to analyze how language models (not necessarily the LLMs) complete various tasks. Standard methods for this analysis include visualizing attention patterns to examine computational properties (such as locality and time invariance) and employing linear probing on the hidden states to determine the extent of learning. Given that the data is synthetic and the ground-truth mappings are generally known, it becomes feasible to isolate the influence of various factors (e.g., prompting strategies, chain-of-thought reasoning, and data formatting). For example, a dataset designed for learning group operations, as detailed in [ZBB+22], facilitates the exploration of how pretraining, data composition, and neural architecture influence reasoning tasks within LLMs. Similarly, the generation of synthetic context-free grammar (CFG) data, as described in [AZL23a], enables training GPT-2 models, uncovering their capacity to learn dynamic programming algorithms for parsing CFGs. Synthetic datasets focusing on biographical knowledge, presented in [AZL23b, AZL23c, AZL24], probe into the mechanisms of knowledge storage, retrieval, manipulation, and the implications of scaling laws. Moreover, the work in [LSL+23] introduces synthetic datasets aimed at understanding how smaller LLMs tackle basic arithmetic operations, like addition, and examines the effects of few-shot prompting, pretraining, and model scaling [LSL+23]. This paper builds upon these investigations by conducting controlled experiments with a path planning dataset, thereby shedding light on the complexities and challenges of planning in LLMs.
8 Discussion and Conclusion
In this paper, we present our investigation on how the Transformer architecture executes the path-finding task, which abstracts a number of planning tasks one may encounter in practical scenarios. We first show that the Transformer has the expressive power of path finding by manually construct a 1-layer 1-head Transformer that encodes the adjacency matrix and the reachability matrix with its weight matrices. Next we provide a theoretical analysis demonstrating that the Transformer can indeed learn the adjacency matrix and the observed reachability matrix (a limited form of reachability) from the path training data through the common gradient descent mechanism on cross-entropy loss function. We then verify our theoretical analysis through comprehensive experiments, both on synthetical network data and on a concrete Blocksworld planning benchmark. Our empirical findings support our theoretical analysis and demonstrate that the Transformer can achieve high accuracy in many cases, it can concentrate its attention correctly to the target node, and it extracts adjacency matrix and observed reachability matrix as predicted by our theoretical analysis. Our theoretical analysis further points out a limitation of the Transformer — it cannot learn well the transitive reachability, and this point is validated by our experimental results.
Our study is the first to combine theoretical analysis with empirical evaluation on such planning tasks. In particular, we are the first to apply theoretical analysis using the gradient descent learning method of the Transformer on a concrete task and provide theoretical interpretation, which is validated by our experimental results. More importantly, our theoretical analysis is able to unveil a potential limitation of the Transformer in completing the path-finding task, which guides us to verify that indeed this limitation occurs in the experiments. This demonstrates the power of theoretical analysis as a guidance for understanding practical Transformers and LLMs.
Although the path-finding task is a simple task on a network and our investigation is still the first attempt to understand Transformer’s capability on this task, our findings may already provide some insights and implications to path finding and planning in general. We can see that, due to the autoregressive nature of learning, the Transformer cannot execute path finding by a simple trial-and-error search such as depth-first or breadth-first search, and thus it has to plan ahead well in generating the next node on the path correctly. We demonstrate that the Transformer can adjust its effort accordingly by concentrating its attention to the target, and generating the next node that is both adjacent to the current node and reachable to the target, mimicking the human intelligence. It suggests that for many planning tasks in general, the Transformer architecture is adjusting its learning by balancing the consideration between the immediate continuation of the next step and its final goal of reaching the target.
Our investigation opens up many possible future directions, and we list several of them below. (a) Extend our study to hyper-graphs and hyper-paths, where a hyper-edge models the situation in which several preconditions need to be satisfied together in order to carry out the next step, which is common in task planning and mathematical proofs. (b) Further understand the limitation of the Transformer on path finding: the current inability of transitive reachability may be due to the limited expressive power of the path language, and it may also suggest the fundamental difference between Transformer’s interpretation and human’s interpretation on path finding. This requires further study, perhaps by enriching the languages, or by an investigation on the deployed LLMs such as GPT4. It can also stimulates new improvement to the Transformer architecture to achieve better performance. (c) Study the connection between the abstract path-finding task and some concrete planning tasks (e.g. block moving in Blocksworld, etc.), to see if Transformer has any capability of abstracting a concrete task to the more abstract path-finding task, or find if what Transformer does for concrete planning tasks has any commonalities with the abstract path-planning task. (d) Study deployed LLMs or fine-tune LLMs to connect their planning capabilities with our path-finding investigation.
We are planning to deepen our investigation on the above fronts. We also hope that our work could inspire more people to study LLMs with combined theoretical and empirical analysis, with the ultimate goal of understanding and explaining the power of LLMs.

### v3

@@ Appendix F Path-planning in Blocksworld
To further validate the theoretical results in Section 3 and the practicability of the proposed path-finding task, we consider Blocksworld benchmark [22]. Blocksworld is a scenario consisting of a set of blocks identified by different colors. The blocks are either placed on table or on top of another block and the task is to plan a block manipulation from the source state to the target state.
@@
We formulate Blocksworld as a path-finding task. Here we construct a graph $G_{BW}$ for the case with $4$ blocks, where each node represents a state of the blocks. For example, node 00 refers to the state that “the red block is on top of the blue block, the blue block is on top of the orange block, the orange block is on top of the yellow block, and the yellow block is on the table”. $G_{BW}$ is a directed graph with $73$ nodes, and the adjacency matrix of $G_{BW}$ is presented in Figure 11(a).
@@
In the original Blocksworld task, the answer is a sequence of actions, which is equivalent to the notion of edges in $G_{BW}$. We reformulated it to let the model output a path from the given source state to the given target state, only consisting of the nodes. This can be seen as a simplified version and a pure planning task. We randomly select $80\%$ of all node pairs for training and the rest $20\%$ for testing, and generate $50000$ training sequences in the same format as introduced in Section 2. We mainly use Transformers with $1$ layer and $1$ head for the convenience of visualization.
@@
Figure 12: Experiment for reachability on Blocksworld benchmark.
F.1 Results
@@
We first present the accuracy results during training when using different embedding sizes $d\in\{30,60,120\}$. As shown in Figure 11(d), although a smaller embedding size results in a longer time to converge, all models reach an accuracy near $100\%$ at the end of the training.
@@
Then, we use the same method introduced in Section 4 to visualize the attention map and the $\bm{W}^{M^{\prime}}$ matrix for the model with $d=120$ after the entire iterations. In Figure 11(c), we can see that when predicting the tokens on the path, almost all the attention weights are on the second token which represents for the target node, demonstrating the capability of the model to learn a correct attention. For adjacency matrix, we find that the $\bm{W}^{M^{\prime}}$ matrix in Figure 11(b) almost perfectly aligns to the real adjacency matrix of $G_{BW}$. And the weight gap (average edge weight minus average non-edge weight) for all models keeps increasing in the training process until convergence, as shown in Figure 11(e).
@@
In addition, we present the results related to observed reachability matrix in Figure 12. Figure 12(a) shows the observed reachability in the training dataset. Although $G_{BW}$ is fully-connected, some reachability are not observed since we request that all training data has no cycle. Specifically, in this case, each of the first $24$ nodes is not observed reachable to any nodes other than itself. To validate whether the Transformer has captured this information, we construct $\bm{W}^{V^{\prime}}$ matrix through the same method presented in Section 4. As shown in Figure 12(b), the first 24 columns of the $\bm{W}^{V^{\prime}}$ matrix are noticeably darker, which aligns with the observed reachability matrix in Figure 12(a). Furthermore, we plot the gap between the average weight of $\bm{W}^{V^{\prime}}$ on observed reachability and the average weight of $\bm{W}^{V^{\prime}}$ on non-observed reachability in Figure 12(c), and find that this gap keeps increasing for all models. Since there does not exist any test pairs with degree 2 or more (as defined in Section 4), we do not compare the accuracy between different degrees in Blocksworld.
@@
In summary, our experimental results on the Blocksworld benchmark confirm our theoretical analyses (Theorem 3) and empirical results on the synthetic data, and they at least partially explain the planning capability of the Transformer on the Blocksworld scenario.